OpenStudy (anonymous):

Simplify i^38 a. i b. -1 c. -i d. 1

OpenStudy (anonymous):

no

OpenStudy (solomonzelman):

I is a square root of -1.

OpenStudy (anonymous):

is the answer 1 thats what i got

OpenStudy (solomonzelman):



OpenStudy (anonymous):

i=i i^2= -1 i^3=-i i^4=1 that repeats

OpenStudy (solomonzelman):

No, it is not 1.

OpenStudy (solomonzelman):

I'll give you tips.

OpenStudy (anonymous):

http://openstudy.com/study#/updates/4e2f42cf0b8ba7b2da4007c6 yeah but ^^

OpenStudy (solomonzelman):